# Dinkel Yan Wang Wang Sun Niu Liu Li Zhang Luan MiLM Plus, Xiaomi Inc.China

### v1

@@ Abstract
keywords:
contrastive language-audio pretraining, general pretraining, general audio encoders, large-language models
@@ 1 Introduction
In the field of computer vision, Contrastive Language-Image Pretraining (CLIP) [1] represents a significant breakthrough in extracting efficient representations that can be applied across various downstream tasks and domains. Similarly, Contrastive Language-Audio Pretraining (CLAP) [2, 3, 4] bridges text and audio domains, enabling zero-shot transfer learning i.e., testing the model on novel concepts that is has not seen during training. Notably, [5] trained on 4.6 Million pairs of audio and speech data, but has shown poor results for trivial speech classification tasks such as keyword spotting (see MSCLAP-2023 in Figure 1). While multilingual extensions [6] improved retrieval performance across eight languages, their approach still lacks basic speech understanding. CLAP embeddings primarily target sound and music, missing comprehensive speech representation (i.e., spoken language) - a critical aspect of audio processing. While there has been previous work focusing on speech-text embeddings using contrastive learning [3], a general approach that can be used between sound, music and speech domains is still missing. This work proposes general language audio pretraining (GLAP), an extension of previous CLAP works, aimed at aligning speech content with text, without compromising in sound and music performance. Our experiments demonstrate that GLAP achieves competitive performance in music and sound retrieval tasks while significantly improving speech understanding capabilities. GLAP also effectively generalizes its speech and sound understanding capabilities beyond English.
Figure 1: GLAP’s retrieval and zero-shot performance. A@T and T@A represent retrieval tasks of Audio-to-Text and Text-to-Audio, respectively, others are zero-shot (number of labels in brackets). Missing baselines were evaluated by the authors.
@@ 2 General language audio pretraining
To enable speech understanding in CLAP models, which are trained on sound and music data, one simple solution is to add speech data to the training dataset. However, as our analysis in Section 4.1 shows, this approach leads to compromised performance due to the absence of a unified audio encoder. The model either performs well on sound/music or on speech, but struggles to excel at both simultaneously. Thus, GLAP has two primary goals:
  1. 1.
    Deliver a unified encoder framework that maintains high performance for sound, music and speech retrieval tasks to enable alignment across these audio modalities.
  2. 2.
    Enable multilingual search capabilities for sound, music and speech content.
GLAP is trained with pairs of audio-text samples $(a,t)$ in contrastive fashion. Features from these audio-text pairs are extracted through a pre-trained multi-lingual text encoder ${\text{E}_{T}}$ and a pre-trained general audio encoder ${\text{E}_{A}}$:
@@
A trainable multi-layer perceptron (MLP) is added to align the dimensions. Finally, the pair $(e_{a},e_{t})$ is scored using cosine distance $s=\frac{e_{a}\cdot e_{t}^{T}}{||e_{a}||\cdot||e_{t}||}$. Unlike previous works, we use the sigmoid loss [7] as our main training objective $\mathcal{L}$, computed as:
@@
Figure 2: The GLAP framework. GLAP is trained via contrastive learning where positive pairs ($\psi[i,j]=1$) are shown in blue and negative pairs ($\psi[i,j]=-1$) in white, with the added benefits of enabling multilingual speech-content retrieval, ontop of the standard sound/music capabilities.
@@ 3 Experimental Setup
The audio data is preprocessed by resampling all datasets to a single channel at 16 kHz. The trainable loss parameters $\tau,\beta$ (Equation 2) are initialized as $0.07$ and $-10$ respectively. For text embeddings, we use the text encoder from Sonar [8] as the default model, following [6]. We use a batch size of 128 per GPU across eight A800 GPUs, resulting in an effective batch size of $B=1024$. Embeddings are gathered across all GPUs before calculating the loss. Each epoch is defined as processing 10,000 batches, with training running for a maximum of 20 epochs. Model training employs an 8-bit Adam optimizer with a cosine decay scheduler. The learning rate starts at 0, warms up to ${10^{-4}}$ over the first two epochs, and decays to ${10^{-5}}$ over the remaining training period. Training takes approximately 1.5 days, with the best models typically achieved within the first 15 epochs. The source code and checkpoints are publicly available111github.com/xiaomi-research/dasheng-glap.
3.1 Datasets
Training
In this work, we integrate a wide range of existing audio-text datasets, as outlined in Table 1. Our primary speech training dataset is YODAS [9], a 400k-hour YouTube corpus labeled mainly via automated speech-to-text pipelines. Due to its noisy labeling, we supplement it with cleaner English (GigaSpeech, LibriSpeech) and Chinese (AISHELL-1) datasets. While YODAS covers 145 languages, relying solely on multilingual speech data limits generalization in sound/music retrieval. To address this, we followed [6] and leveraged Sonar [8] to translate the original English captions of all sound and music datasets into other seven widely spoken languages: German, Chinese, Catalan, Spanish, Japanese, French, and Dutch. As shown in Table 1, the training data is skewed towards speech, leading to poor performance on non-speech tasks. To balance this, we categorize the data into four groups: sound + music, English speech (GigaSpeech + LibriSpeech + YODAS English), Chinese speech (AISHELL-1 + YODAS Chinese), and other languages in YODAS. During training, we sample equally from each group, ensuring a balanced training process.
@@ 3.1 Datasets
Table 1: Training datasets with duration (hours), audio-text pairs (# Pairs), and languages (# Lang). Music and sound data, labeled in English, are auto-translated into seven other languages via Sonar.
@@
Table 2: Retrieval results for music and speech datasets. indicates evaluation from a public checkpoint. Best in bold; higher is better.
@@
We evaluate retrieval performance across seven test sets, including sound datasets such as Auto-ACD (ACD)[14], AudioCaps (AC)[17], Clothov2 (Clotho) [18], music datasets such as MusicCaps (MC) [19] and the speech datasets LibriSpeech (LS) [11], GigaSpeech [10] and AISHELL-2 (AIS2) [23]. For zero-shot classification, we primarily follow [5] and evaluate on ESC-50, FSD50K, UrbanSound8K (US8K), CREMA-D, GTZAN, NSynth instruments, Beijing-Opera, VocalSound, as well as Speech Commands V1/V2 (SCV1/2) [24] and Fluent Speech Commands (FSC) [25]. All test datasets with the exception of AIS2 are labeled in English.
3.2 Evaluation metrics
In audio-text retrieval tasks, performance is evaluated using recall at rank (R@k), where R@k is 1 if the target item appears in the top k retrieved items, otherwise 0. We also use mean average precision at rank 10 (mAP10) for a more comprehensive comparison. For zero-shot inference, accuracy is used for single-class classification, and mean average precision (mAP) is used for multi-label classification, as is standard practice [5].
3.3 Prompting
GLAP supports zero-shot inference, allowing the model to generate outputs directly from text prompts without prior training on specific tasks. In zero-shot scenarios, crafting an effective text prompt is crucial for achieving optimal performance. This is particularly important for GLAP, as it lacks a dedicated token to distinguish between a spoken word (e.g., “cat”) and a sound event (e.g., “the sound of a cat”). Prompts used in this work are depicted in Table 3.
@@ 4.1 Audioencoder investigation
Given the extensive range of previously explored CLAP methods, one might question why earlier approaches failed to achieve strong performance in general audio-text pretraining. In our view, a key limitation lies in the reliance on sound-event audio encoders for CLAP, which we believe represents a significant bottleneck for performance. In this section, we examine the role of audio encoders in general contrastive audio-text learning. For this purpose, we compare the proposed training framework using five different audio-encoders, being Dasheng [26], CED-Base [27], Beats [28], Whisper-Base [29] and WavLM [30]. Each encoder is selected for its approximately 90M parameters and support for variable-length inputs, a crucial requirement for processing speech recognition datasets—a capability lacking in many CLAP audio encoders [2, 5, 3]. Each encoder is initialized from a publicly available checkpoint and trained on the same dataset described in Section 3.1, using the previously described settings (Section 3), with Sonar serving as the default text encoder.
As it can be seen in our results in Table 4, the choice of audio-encoder is vital to achieve a well-balanced performance across tasks. Sound-event encoders, such as CED and Beats, perform well in sound and music retrieval tasks but struggle with speech-related tasks. Conversely, Whisper and WavLM excel in speech-related retrieval but underperform in sound event and music datasets. Dasheng, on the other hand, proves to be the most versatile choice for general audio encoding, achieving competitive performance across sound, music, and speech domains. Based on these findings, all subsequent experiments utilize a single GLAP model with Dasheng as its audio encoder, without fine-tuning on target datasets.
@@
Table 4: Text-to-Audio retrieval performance across five datasets, categorized by domain. LS-other refers to the test-other subset of LibriSpeech, while AIS2 corresponds to the AISHELL-2 test set. All experiments were conducted using the proposed training dataset and configuration. Values indicate mAP10, where higher scores represent better performance.
Method AudioCaps Clotho
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
BLAT [4] 33.3 82.4 40.4 85.7 12.3 46.1 13.9 48.2
LClap-Large [3] 34.2 84.1 43.1 90.1 15.3 51.2 20.8 60.0
MSCLAP-2022 [2] 33.5 80.2 47.8 90.7 16.2 51.4 23.6 60.3
MSCLAP2023 [5] 35.6 - 42.5 - - 15.7 22.9 -
Wavcaps-CNN14 [16] 34.7 82.5 44.7 86.2 21.2 59.4 25.9 65.8
Wavcaps-HTSAT [16] 39.7 86.1 51.7 90.6 20.2 58.8 26.5 67.3
Auto-ACD [14] 39.5 85.4 53.7 91.7 15.3 52.1 17.7 52.6
T-CLAP [31] 39.7 86.9 49.8 91.9 17.3 53.6 21.8 57.4
MLCLAP [6] 40.7 87.8 50.1 92.8 18.8 59.0 21.1 62.5
Cacophony [32] 41.0 86.4 55.3 92.4 20.2 58.8 26.5 67.3
SoundVECaps [13] 41.2 85.3 53.3 93.0 - - - -
GLAP 41.7 86.1 54.4 91.1 19.4 58.3 21.8 61.5
Table 5: Sound event retrieval results compared to baselines.
@@ 4.3 Music and speech retrieval results
For speech retrieval, we select the test-other dataset from Librispeech as a representative in-domain English speech benchmark, while the AISHELL-2 test set serves as an unseen Chinese evaluation set. To assess music-related capabilities, we use the widely adopted MusicCaps dataset [19]. Since some baseline models operate at different sampling rates (16 kHz vs. 44/48 kHz), we resample the test datasets accordingly. As shown in Table 2, our approach significantly outperforms previous methods in both music and speech retrieval. GLAP demonstrates exceptional multilingual retrieval performance, achieving over 93% on the English LibriSpeech-other test set and 98% on the Chinese AISHELL-2 test set. While most compared CLAP models offer a competitive performance on the MC dataset.
@@ 4.4 Zero shot evaluation
In this section, the zero-shot capabilities our approach are assessed using the prompts provided in Table 3. True zero-shot evaluation is difficult to determine, as some studies are trained on the respective datasets [3, 5] using its labels, meaning they are not strictly zero-shot. As shown in Table 6, our approach performs similarly to other baselines in sound-event and music classification. However, it excels in the keyword spotting task, significantly outperforming the baselines with accuracies of 95.8%, 96.6%, and 75.6% for the SCV1, V2, and FSC datasets, respectively. Notably, the FSC dataset requires understanding entire sentences, not just individual words, highlighting our model’s ability to align well with spoken content, as well as detecting single-word keywords (SCV1/2).
@@ 4.5 Multilingual capabilities
GLAP’s multilingual spoken content capabilities are assessed through a zero-shot evaluation on the Multilingual Spoken Words (MSW) Corpus [33].
Figure 3: Multilingual zero-shot keyword spotting (KWS) performance across 50 languages. Only the test set for each language is used, and the accuracies are reported. The number of keywords (num) for each language is shown on the right.
Results seen in Figure 3 show a strong zero-shot capability of GLAP on fifty languages. Oriya and Guarani achieve the best performance at 70% and 65.9% respectively. Notably, Chinese achieves 57.4% accuracy with 496 keywords, while Russian, containing 15,844 keywords, results in 39.6%. Additionally, we assess GLAP’s multilingual sound and music understanding by conducting zero-shot evaluations on the ESC-50, US8K and GTZAN datasets. For this, we use ChatGPT to translate the original (English) labels into the target language and adjust the prompts (Table 3) accordingly. As shown in Table 7, while performance drops compared to the English baseline, GLAP remains effective in multilingual sound-event classification. The model performs impressively in Russian (Ru) despite being trained only on Russian speech-text pairs from YODAS, not music/speech pairs (see Section 3.1). This suggests effective transfer of multilingual text-based knowledge to the audio domain.
@@ 5 Conclusion
We introduce GLAP, a versatile language-audio pretraining framework that enables multilingual and multi-domain modeling of both audio and text. To the best of our knowledge, it is the first single system to integrate general audio and text embeddings into a unified contrastive framework. GLAP demonstrates competitive performance on well-established benchmarks like AudioCaps and Clotho, while surpassing previous methods in music and speech retrieval tasks. Zero-shot evaluations show strong results for English sound and music tasks, extending effectively to other languages. Inference on the Multilingual Spoken Words dataset highlights robust multilingual capabilities beyond English.

### v2

# GLAP: General contrastive audio-text pretraining across domains and languages

@@ Abstract
Index Terms—  contrastive language-audio pretraining, general pretraining, general audio encoders, large-language models
@@ 1 Introduction
In the field of computer vision, Contrastive Language-Image Pretraining (CLIP) [21] represents a significant breakthrough in extracting efficient representations that can be applied across various downstream tasks and domains. Similarly, Contrastive Language-Audio Pretraining (CLAP) [25, 26] bridges text and audio, enabling zero-shot transfer learning. However, existing CLAP models, including those trained on large datasets like the one in [12] and multilingual extensions [27], perform poorly on basic speech tasks like keyword spotting. CLAP embeddings primarily target sound and music, missing comprehensive speech representation (i.e., spoken language) - a critical aspect of audio processing. This work proposes general language audio pretraining (GLAP) aimed at aligning speech content with text, without compromising in sound and music performance. GLAP further effectively generalizes its speech and sound understanding capabilities beyond English which be seen in Figure 1.
Fig. 1: GLAP’s retrieval and zero-shot performance. A@T and T@A represent retrieval tasks of Audio-to-Text and Text-to-Audio, respectively, others are zero-shot (#labels).
@@ 2 General language audio pretraining
To enable speech understanding in CLAP models, which are trained on sound and music data, one simple solution is to add speech data to the training dataset. However, as our analysis in Section 4.1 shows, this approach leads to compromised performance due to the absence of a unified audio encoder. Thus GLAP’s main goal is to provide a unified framework framework that maintains high performance for sound, music and speech retrieval tasks, enabling alignment across these audio modalities. GLAP is trained with pairs of audio-text samples $(a,t)$ in contrastive fashion. Features from these audio-text pairs are extracted through a pre-trained multi-lingual text encoder $\text{E}_{T}$} and a pre-trained general audio encoder $\text{E}_{A}$}:
@@
[…]
@@
Fig. 2: GLAP enables multilingual speech-content retrieval, ontop of the standard sound/music capabilities.
@@ 3 Experimental Setup
All audio data is preprocessed by resampling all datasets to a single channel at 16 kHz. Following [27], we use the text encoder from Sonar [10] for text embeddings. We initialize the loss parameters as $\tau=0.07,\beta=-10$ (Equation 2). We train the model with an effective batch size of $B=1024$ (128 per GPU) for a maximum of 20 epochs (10,000 batches per epoch). Training uses an 8-bit Adam optimizer with a cosine decay scheduler, where the learning rate warms up from 0 to $10^{-4}$} over the first two epochs and then decays to $10^{-5}$}. The source code and checkpoints are publicly available111github.com/xiaomi-research/dasheng-glap.
@@
Table 1: Training datasets with duration (hours), audio-text pairs (# Pairs), and languages (# Lang). Music and sound data labeled in English were auto-translated to 7 additional languages via Sonar.
@@
Table 2: Music/Speech retrieval results. indicates evaluation from a public checkpoint. Best in bold; higher is better.
3.1 Datasets
Training
We trained our model on a diverse range of audio-text datasets, as detailed in Table 1. Our main dataset is YODAS [14], a 400k-hour YouTube corpus primarily labeled with noisy automated speech-to-text pipelines. To improve data quality, we supplemented YODAS with cleaner English datasets (GigaSpeech, LibriSpeech) and a Chinese dataset (AISHELL-1). While YODAS covers 145 languages, relying solely on multilingual speech data limits generalization in sound/music retrieval. To address this, we followed [27] and leveraged Sonar [10] to translate the original English captions of all sound and music datasets into other seven widely spoken languages: German, Chinese, Catalan, Spanish, Japanese, French, and Dutch. To address the imbalance of speech data shown in Table 1, we categorized the data into four groups: sound + music, English speech, Chinese speech, and other languages and sample equally from each group.
@@ 3.1 Datasets
We evaluate retrieval performance across sound datasets such as Auto-ACD (ACD) [23], AudioCaps (AC) [13], Clothov2 (Clotho) [9], music datasets such as MusicCaps (MC) [1] and the speech datasets LibriSpeech (LS) [20], GigaSpeech [4] and AISHELL-2 (AIS2) [3]. For zero-shot classification, we primarily follow [12] and evaluate on ESC-50, FSD50K, UrbanSound8K (US8K), CREMA-D, GTZAN, NSynth instruments, Beijing-Opera, VocalSound, as well as Speech Commands V1/V2 (SCV1/2) and Fluent Speech Commands (FSC) [15]. All test datasets with the exception of AIS2 are labeled in English.
3.2 Evaluation metrics and Zero-shot Prompting
Audio-text retrieval performance is measured using recall at rank (R@k), which checks if the correct item is in the top k results. For a more complete comparison, we also use mean average precision at rank 10 (mAP10). For zero-shot inference, we use accuracy for single-class classification and mean average precision (mAP) for multi-label classification, consistent with standard practice [12]. GLAP performs zero-shot inference, but without a dedicated token to differentiate between spoken words and sound events, the model relies on effective text prompts. Prompts used in this work are shown in Table 3.
@@ 4.1 Audioencoder investigation
In our view, a key limitation lies in the reliance on sound-event audio encoders for CLAP. For this purpose, we compare the proposed training framework using five distinct audio-encoders, being Dasheng [8], CED-Base [7], Beats [6], Whisper-Base [22] and WavLM [5], each with $\approx$ 90M parameters and support for variable length inputs. Each encoder is initialized from a publicly available checkpoint and trained using the settings in Section 3. As the results in Table 4 indicate, sound-event encoders like CED and Beats excel at music and sound tasks, but perform poorly on speech. Conversely, Whisper and WavLM are strong on speech but weak on other audio types. Dasheng is the most versatile, performing well across all domains. Because of this, further experiments use Dasheng as the primary audio encoder.
@@
Table 4: Text-to-Audio retrieval performance across five datasets in mAP10, where higher is better. LS-other represents test-other of LibriSpeech and AIS2 represents the AISHELL-2 test set.
@@
Method AudioCaps Clotho
Text-to-Audio Audio-to-Text Text-to-Audio Audio-to-Text
R@1 R@10 R@1 R@10 R@1 R@10 R@1 R@10
BLAT [26] 33.3 82.4 40.4 85.7 12.3 46.1 13.9 48.2
LClap-Large [25] 34.2 84.1 43.1 90.1 15.3 51.2 20.8 60.0
MSCLAP2023 [12] 35.6 - 42.5 - - 15.7 22.9 -
Wavcaps-CNN14 [18] 34.7 82.5 44.7 86.2 21.2 59.4 25.9 65.8
Wavcaps-HTSAT [18] 39.7 86.1 51.7 90.6 20.2 58.8 26.5 67.3
Auto-ACD [23] 39.5 85.4 53.7 91.7 15.3 52.1 17.7 52.6
MLCLAP [27] 40.7 87.8 50.1 92.8 18.8 59.0 21.1 62.5
SoundVECaps [28] 41.2 85.3 53.3 93.0 - - - -
GLAP 41.7 86.1 54.4 91.1 19.4 58.3 21.8 61.5
Table 6: Sound event retrieval results compared to baselines.
@@ 4.3 Music and speech retrieval results
We assessed GLAP’s retrieval capabilities using the Librispeech test-other dataset for English speech and the AISHELL-2 test set for unseen Chinese. For music, we used the MusicCaps dataset [1]. As shown in Table 2, GLAP significantly outperforms previous methods in both music and speech retrieval. It achieves over 93% on the English LibriSpeech test set and 98% on the Chinese AISHELL-2 test set, demonstrating exceptional multilingual performance.
@@ 4.4 Zero shot evaluation
Zero-shot results are shown in Table 5. Our model performs similarly to other baselines for sound and music classification. However, it significantly outperforms them in keyword spotting tasks, achieving 95.8%, 96.6%, and 75.6% accuracy on the SCV1, V2, and FSC datasets, respectively.
@@ 4.5 Multilingual capabilities
GLAP’s multilingual capabilities are evaluated using zero-shot tests on the Multilingual Spoken Words (MSW) Corpus [17].
Fig. 3: Multilingual zero-shot keyword spotting performance across 50 languages. The number of keywords (num) for each language is shown on the right.
Results in Figure 3 show strong performance across 50 languages. For example, Oriya and Guarani achieve the best accuracy at 70% and 65.9%, respectively. Notably, Chinese, with 496 keywords, reaches 57.4% accuracy, while Russian, with 15,844 keywords, achieves 39.6%. We also assessed GLAP’s multilingual sound and music understanding on the ESC-50, US8K, and GTZAN datasets. As shown in Table 7, while performance decreases compared to the English baseline, GLAP remains effective for multilingual sound classification.
@@ 5 Conclusion
We introduced GLAP, a versatile language-audio pretraining framework that enables multilingual and multi-domain modeling of both audio and text. To the best of our knowledge, it is the first single system to integrate general audio and text embeddings into a unified contrastive framework. GLAP demonstrates competitive performance on well-established benchmarks like AudioCaps and Clotho, while surpassing previous methods in music and speech retrieval tasks. Zero-shot evaluations show strong results for English sound and music tasks, extending effectively to other languages. Inference on the Multilingual Spoken Words dataset highlights robust multilingual capabilities beyond English.